OpenStudy (anonymous):

Assume that y varies inversely with x. If y = -21 when x = 4, find x when y = 7. A. x = - 1.3 B. x = -12 C. x = -3

OpenStudy (anonymous):

y=k/x k=yx k=(-21)(4)=-84 y=-84/x x=-84/y x=-84/7=-12

OpenStudy (anonymous):

thnk u

OpenStudy (anonymous):

yw